# On codes induced from Hadamard matrices  [^1] Keywords: Hadamard, code, linear, convolutional, self-dual, dual-containing, quantum code, linear complementary AMS Classification 2020: 94B05, 94B10, 15B99

### v2

@@ 1 Introduction
Higher memory convolutional codes may also be generated by breaking the Hadamard matrices further into blocks ; see [27] for this.The general process for constructing higher memory convolutional codes from Hadamard matrices is left for later development; however an example, 2.10, is given for a small case to show how the process can proceed for Hadamard matrices.
@@ 1.1 Additional notation and background
Convolutional codes can be catastrophic or non-catastrophic; see for example [3] for the basic definitions. A catastrophic convolutional code is prone to catastrophic error propagation and is not much use. A convolutional code described by a generator matrix with right polynomial inverse is a non-catastrophic code; this is sufficient for our purposes. The designs given here for the generator matrices allow for specifying directly the control matrices and the right polynomial inverses. Lack of algebraic construction methods for designing convolutional codes limited their size and availability, see McEliece [3] for discussion and also [10, 11, 21, 34]. It is shown here how Hadamard matrices can be used to construct convolutional codes but see also [27, 28]. Several algorithms exist for decoding convolutional codes, the most common ones being the Viterbi algorithm and the sequential decoding algorithm.
@@ 2.1 Ternary codes from Hadamard matrices
It is interesting to find the distances attained. For a self-dual $n,\frac{n}{2},d]_{3}$ ternary code it is known that $d\leq\lfloor\frac{n}{12}\rfloor+3$ [29]. For $n=12k$ extremal ternary self-dual codes exist for lengths $n=12,24,36,48,60$ and do not exist for $n=72,96,120$ and for $n\geq 144$. Now by Corollary 2.17 $\frac{n}{2}$ linearly independent rows over $GF(3)$ of a Hadamard matrix of size $n=12k$ generate a self-dual ternary code; when $n=12,24,36,48,60$ it is verified by computer that these are optimal.
@@
(i) Suppose $A$ consists of $\frac{n}{2}$ rows of $H$ and $B$ consists of the other $\frac{n}{2}$ rows of $H$. Then $G(z)=A+iB$ with $i=\sqrt{-1}$ in $GF(3^{2})$ generates a self-dual convolutional $(n,\frac{n}{2},\frac{n}{2};1,d)$ code in $GF(3^{2})$ where $d=d(A)+d(B)$.
(ii) Suppose $A$ consists of $r$ rows of $H$ with $r>\frac{n}{2}$ and $B$ consists of the other (nr)(n-r) rows of $H$. Define $B_{1}=\left(\begin{smallmatrix}\underline{0}\\ B\end{smallmatrix}\right)$ where $\underline{0}$ is the zero $(2n-r)\times n$ matrix. Then $G(z)=A+iB_{1}x$ generates a dual-containing convolutional (n,r,nr;1,d)(n,r,n-r;1,d) code in $GF(3^{2})$ where $i=\sqrt{-1}$.

### v3

@@ 1 Introduction
Higher memory convolutional codes may also be generated by breaking the Hadamard matrices further into blocks ; see [26, 25] for this.The general process for constructing higher memory convolutional codes from Hadamard matrices is left for later development; however an example, 2.10, is given for a small case to show how the process can proceed for Hadamard matrices.
@@ 1.1 Additional notation and background
Convolutional codes can be catastrophic or non-catastrophic; see - for example [3] for the basic definitions. A catastrophic convolutional code is prone to catastrophic error propagation and is not much use. A convolutional code described by a generator matrix with right polynomial inverse is a non-catastrophic code; this is sufficient for our purposes. The designs given here for the generator matrices allow for specifying directly the control matrices and the right polynomial inverses. Lack of algebraic construction methods for designing convolutional codes limited their size and availability, see McEliece [3] for discussion and also [10, 11, 20, 32]. It is shown here how Hadamard matrices can be used to construct convolutional codes but see also [26, 27]. Several algorithms exist for decoding convolutional codes, the most common ones being the Viterbi algorithm and the sequential decoding algorithm.
@@ 2.1 Ternary codes from Hadamard matrices
It is interesting to find the distances attained. For a self-dual $n,\frac{n}{2},d]_{3}$ ternary code it is known that $d\leq\lfloor\frac{n}{12}\rfloor+3$ [mallow]. For $n=12k$ extremal ternary self-dual codes exist for lengths $n=12,24,36,48,60$ and do not exist for $n=72,96,120$ and for $n\geq 144$. Now by Corollary 2.17 $\frac{n}{2}$ linearly independent rows over $GF(3)$ of a Hadamard matrix of size $n=12k$ generate a self-dual ternary code; when $n=12,24,36,48,60$ it is verified by computer that these are optimal.
@@
(i) Suppose $A$ consists of $\frac{n}{2}$ rows of $H$ and $B$ consists of the other $\frac{n}{2}$ rows of $H$. Then $G(z)=A+iBz$ with $i=\sqrt{-1}$ in $GF(3^{2})$ generates a self-dual convolutional $(n,\frac{n}{2},\frac{n}{2};1,d)$ code in $GF(3^{2})$ where $d=d(A)+d(B)$.
(ii) Suppose $A$ consists of $r$ rows of $H$ with $r>\frac{n}{2}$ and $B$ consists of the other (nr)(n-r) rows of $H$. Define $B_{1}=\left(\begin{smallmatrix}\underline{0}\\ B\end{smallmatrix}\right)$ where $\underline{0}$ is the zero $(2n-r)\times n$ matrix. Then $G(z)=A+iB_{1}z$ generates a dual-containing convolutional (n,r,nr;1,d)(n,r,n-r;1,d) code in $GF(3^{2})$ where $i=\sqrt{-1}$.